Bolometer Array Development -- Brian

  • work completed/decisions made

  • upcoming tasks/decisions
    • Completed drawings of the new dewar layout with final clearance requirements is expected from the UPenn team next week. Once our team has had time to adequately review the drawings a CDR will be held to finalize the base design so that the NRAO team can do detailed design of the mounting schemes and rotator.
    • Target date for CDR is the week of 6-Jan.

PTCS program -- Marty

  • Servo
    • work completed/decisions made
      • During maintenance this week the following tasks were accomplished in advance of another round of tests during maintenance next week
        1. downloading images for embedded code fixes to all MCI's
        2. Added test jumpers/wires to permit examination of MCI filter BW/phase on closed loop response
        3. Added a break-out to the input of the 2-Axis analog control board to permit a function generator to be substituted for the MRU potentiometer for rate command injection while in MRU mode
        4. Fixed wiring phasing issues with some azimuth tachs
        5. Balancing IRIG signals from the control room through the servo room into the new CCU computer to improve signal reliability and look of the CCU time. This work may continue into next week.
      • Items #3 7 #4 are to do additional tests to characterize the roughness/oscillation reported in last week's meeting.
    • upcoming tasks/decisions
      • Jason Ray will join the kernel development team to adapt the model for the elevation axis, which is nearing completion, to a model for the azimuth axis.
      • Request for additional motion tests next week
  • Surface
    • work completed: new OOF trajectories created for W-band to put the center of the mapped area between the 2 beams.
    • upcoming tasks: analyze the OOF map from Jan 19 to determine if single-beam W-band maps will produce useful surface maps

20m Educational Program - Marty

  • work completed/decisions made
    • The cal diode and frequency switching capabilities are now built into the wideband mode of the 20m/Skynet backend. With only a few more loose ends tied up, the backend will be ready to turn over to UNC for their interface work. The narrowband work will be suspended until which time the VEGAS narrowband modes are complete and can be re-purposed for the 20m Spectrometer.
    • The 20m project management met and assigned the remaining tasks for the X-Band cryo refurbishment, the L-Band rehabilitation, and backend development to cover for the upcoming loss of Gary and Srikanth from the 20m team.
    • Other discussions included the overall integration of the components in the recorder room, the UNC milestones and need to clarify some of the project responsibilities, and the need to carefully balance the educational aspects of the project with the desire to deliver some capabilities in the near future.
    • Also discussed were the switching signals and LO control, and need to review the configuration and assign tasks before the next meeting.

  • upcoming tasks/decisions
    • There is still some concern by the GB team that the UNC team will face quite a bit of difficulty in dealing with the data from the 20m spectrometer. To help mitigate this risk, the leader of the UNC software team will be invited to join the bi-weekly 20m/Skynet meetings.
    • A proposal is in the works for a summer intern to integrate the narrowband modes from VEGAS into the 20m system and make any other changes that come up over the course of integration with UNC.

-- MartyBloss - 2012-01-26
Topic revision: r1 - 2012-01-26, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback